OpenStudy (anonymous):

(0/5) divided by (3/5) [show steps please]

OpenStudy (anonymous):

(0/5) / 3/5 = (0/5)(5/3) the fives cancel each other out.. so it's 0/3... which is zero

Directrix (directrix):

0/5 / (3/5) = 0 * 5/3 = 0

OpenStudy (anonymous):

Thank you, everyone! c: